OpenStudy (anonymous):

A square pyramid has side lengths of 10 units and a height of 7.6 units. The slant height of each lateral face is 9.1 units. What is the surface area of the pyramid to the nearest unit?

OpenStudy (kainui):

With all my math and physics problems, I like to draw them out so I don't have to keep things in my head and confuse myself. Then I can draw on it and write in things I don't know on top of it. Here we go...  You can see that big H is the height of the pyramid while little h is the height of the slanted side and S is the length of a side. So what information do we need to find out from this?

OpenStudy (anonymous):

 See the previous drawing by kainui... S, h are the base length and slant height respectively.. Perpendicular ht. H is not needed... Area= 10^2 + 2 * 10 * 9.1 = 100+182= 282